I need to solve for x
25=5*10^(x/2)
I know I am supposed to use log to bring the x/2 down somehow but I just can't figure it out

Answer by Alan3354(21583)

I need to solve for x
25=5*10^(x/2)
Divide by 5
5 = 10^(x/2)
--------------
No problem so far, right?
Now take the log of both sides. Logs (base 10) are the exponents to which 10 is raised to get the number. Since the right side is already 10 to an exponent, the exponent is the log.
That is, the log of 10^anything is...anything. In this case, it's x/2.
-------------
log(5) = x/2
x = 2*log(5)
x = log(25)
x =~ 1.39794